OpenStudy (anonymous):

given u=(0,4,3)^T a unit vector that is perpendicular to u is...

OpenStudy (anonymous):

@satellite73 plz help

OpenStudy (anonymous):

@thomaster i really need help

OpenStudy (anonymous):

what is ^T ?

OpenStudy (anonymous):

transpose

OpenStudy (perl):

he means a column vector , since it is a row right now

OpenStudy (perl):

there are infinitely many perpindicular vectors

OpenStudy (anonymous):

hehe jus testing u :) how would u find that?

OpenStudy (perl):

a vector A is perpindicular to vector B when A dot B = 0. so we can first find a perpindicular vector (0,4,3) . (a,b,c) = 0 0a + 4b + 3c = 0 4b + 3c= 0. set a = 1 , b =-3 . then we have 4(-3) + 3c = 0 3c = 12 c = 4. so a perpindicular vector is <1,-3,4> . and you can normalise that by dividing by length of the vector. sqrt( 1^2 + (-3)^2 + 4^2 ) = sqrt( 1 + 9 + 16 ) More generally we have a = a ,b=b , c = -4b/3 so a perpindicular vector is (a, b, -4b/3) = a(1 , 0 , 0 ) + b ( 0, 1, -4/3) . now normalise that

OpenStudy (anonymous):

bit different explanation: (0,4,3) . (a,b,c) = 0 0a + 4b + 3c = 0 for the last expretion to be 0, a can be anything and: 4b=-3c so c=-4/3b What this means? It says, take a R^3 vector in the form (a,b,c) where a is anything, b is anything and c=-4/3b and it will be perpendicular to u

OpenStudy (perl):

but you need a unit vector, so you have to normalize that

OpenStudy (anonymous):

ya, that's the last step

OpenStudy (anonymous):

to write it all: \(v=\frac{(a,b,-4/3b)}{\sqrt{a^2+b^2+(16/9)b^2}}\)

OpenStudy (anonymous):

thank guys